OpenStudy (lolipop_drum312):

someone help Solve for x. −32>−5+9x

pooja195 (pooja195):

Any ideas on where to start?

OpenStudy (lolipop_drum312):

mmm i think its suppose to be < or add and subtract somethin first

OpenStudy (lolipop_drum312):

idk.....

pooja195 (pooja195):

you're on the right track :D any ideas on what we should add?

OpenStudy (lolipop_drum312):

mmmm 5 and 9

pooja195 (pooja195):

Nope they are not like terms what would we add 5 to, to cancel it out so that we have 9x ALONE

OpenStudy (lolipop_drum312):

mmm add it to 32 or 5 itself

pooja195 (pooja195):

good -32+5=?

OpenStudy (lolipop_drum312):

38

OpenStudy (lolipop_drum312):

then we subtract or divide?

pooja195 (pooja195):

That's incorrect -32+5 now 32+5

pooja195 (pooja195):

*not

OpenStudy (nurali):

1. add 5 both side 2. divide by 9 both side

OpenStudy (1davey29):

@Nurali has it right. You would solve it like a normal "solve for x" problem.

OpenStudy (lolipop_drum312):

sorry lagged....so that would be 9 divided by 10 equals 1?

OpenStudy (lolipop_drum312):

so the answer is 1

OpenStudy (lolipop_drum312):

?

pooja195 (pooja195):

Nope tell me what -32+5=?

OpenStudy (lolipop_drum312):

27?

pooja195 (pooja195):

almost! -27

pooja195 (pooja195):

divide both sides by 9 now 9/9 cancels out -27/9=?

OpenStudy (lolipop_drum312):

18

pooja195 (pooja195):

nope we are dividing

OpenStudy (lolipop_drum312):

oooh

OpenStudy (lolipop_drum312):

3

pooja195 (pooja195):

a negative divided by a positive is a negative

pooja195 (pooja195):

So what is our final answer?

OpenStudy (lolipop_drum312):

-3?

pooja195 (pooja195):

Yep! :)

pooja195 (pooja195):

But how would we put it in answer form

OpenStudy (lolipop_drum312):

oh shoot i'm terrible at that

pooja195 (pooja195):

___> x Insert -3 in the blank :P

OpenStudy (lolipop_drum312):

oh ok

OpenStudy (lolipop_drum312):

thank you

pooja195 (pooja195):

yw :)